write an equation in slopw-intercept form if possible of the line trough the pair of points (-7,9) and (-4,5)
Answer by tutorcecilia(2152)

First, I would find the slope of the line:
m = (y2-y1)/(x2-x1) = (5-9)/(-4-(-7))= -4/3
.
Use the slope-intercept formula (y=mx+b) to find the y-intercept(b).
Pick point (-7,9) and plug-in the values:
(9)=(-4/3)(-7)+ b
9 = 28/3 + b
9-28/3 = b
9/1-28/3=b
-1/3 = b
.
Plug the slope (m) and the y=intercept (b) into the formula y-mx+b:
y=(-4/3)x -1/3
.
Check the formula by plugging in one of the given points. Pick point (-7,9):
(9)=(-4/3)(-7) - 1/3
9=9 Checks out.